Errata to David Wells’ masterpiece “Games and Mathematics” Subtle Connections

Doron ZEILBERGER^{1}

I thorougly enjoyed this beautiful book! Here are some minor errors and typos. p.23, 6 up: “96 solutions” should be “92 solutions” p. 50, 7 down: 82 should be 81 p. 71, middle, the power 1/4 should be 1/2 and the upper limit of the second product should be n (and possibly other fixes) p. 109, line 7 up: x = π/4 should be x = 1; line 4 up x = π/6 should be x = 1/\sqrt{(}3) ; line 3 up 1/12 - should be 1/\sqrt{(}3) p. 147, line 9 down: BPP formula is wrongly stated ( lower limit of ∑ is i = 1, 8i − 6 should be 8i + 6) p. 191, two last lines should be deleted p. 215: π → h in various places

1 Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen

Rd., Piscataway, NJ 08854-8019, USA. zeilberg at math dot rutgers dot edu , http://www.math.rutgers.edu/~zeilberg/ . Feb. 19, 2013 Accompanied by Maple package XXX